3. You discover an annual invasive plant (which has discrete reproduction and dies in one year)in a previously un-infested ditch near your house. Upon further investigation, you determine that there are exactly 500 individuals in the population. If this population has a per capita birth rate of 100 plants per year, what is λ for this population (include 3 decimal places)?
Answer #1

Lambda= ln(Nt+1/Nt)

Here, Nt+1= 600 (Since the birth rate= 100/yr)

= ln (600/500) = ln (1.2)= 0.182